# Learning to Play Against Unknown Opponents

### v1

@@ Abstract
We consider the problem of a learning agent who has to repeatedly play a general sum game against a strategic opponent who acts to maximize their own payoff by optimally responding against the learner’s algorithm. The learning agent knows their own payoff function, but is uncertain about the payoff of their opponent (knowing only that it is drawn from some distribution $\mathcal{D}$). What learning algorithm should the agent run in order to maximize their own total utility?
We demonstrate how to construct an $\varepsilon$-optimal learning algorithm (obtaining average utility within $\varepsilon$ of the optimal utility) for this problem in time polynomial in the size of the input and $1/\varepsilon$ when either the size of the game or the support of $\mathcal{D}$ is constant. When the learning algorithm is further constrained to be a no-regret algorithm, we demonstrate how to efficiently construct an optimal learning algorithm (asymptotically achieving the optimal utility) in polynomial time, independent of any other assumptions. Both results make use of recently developed machinery that converts the analysis of learning algorithms to the study of the class of corresponding geometric objects known as menus.
@@ 1.1 Our Results
    Finally, we provide some evidence to show that the unconstrained problem above is hard when both $mn$ and $k$ are large. In particular, we prove that it is NP-hard to take as input a $k$-tuple $(u_{1},u_{2},\dots,u_{k})$ of real numbers and decide whether there exists a learning algorithm $\mathcal{A}$ where the optimizer of type $i$ receives utility at most $u_{i}$ when best responding to $\mathcal{A}$ (Theorem 4.15). This operation is a key subproblem in both of our approaches above, and implies that any efficient construction of an optimal learning algorithm must take a substantially different approach.
@@
Finally, we conclude by emphasizing that although we have stated the above results in the context of maximizing the expected utility of the learner when the optimizer is drawn from a known prior distribution, this optimization framework is fairly generic and allows for optimizing many other objectives over the space of learning algorithms (essentially, any objective which can be written as a concave function of the CSP assignment). Notably, the same techniques allow us to solve the analogous problem in the setting of robust optimization, where the learner instead wants to minimize their worst-case utility against any optimizer in the support of $\mathcal{D}$.
@@ 1.2 Related Work
There is also a rich line of work focused on the problem of learning the (single-shot) Stackelberg equilibrium of an unknown game via repeated interaction with an agent of some specified behavior. For example, one can try to estimate this Stackelberg equilibrium by repeatedly interacting with an agent that myopically best-responds every round [24, 3, 31], a no-regret agent [7, 18], or a no-swap-regret agent / an agent that best-responds to calibrated action forecasts [7, 21]. It is worth emphasizing that this is a fairly different computational problem than the problem we study in this work – although the algorithms we commit to may “try” to learn the payoffs of their opponent in a similar way to these algorithms, we 1. do not make any behavioral assumptions about the opponents (allowing them to react fully rationally and non-myopically), and 2. attempt to find the Bayesian Stackelberg equilibrium of the full repeated game, not the single-shot Stackelberg equilibrium. Indeed, the recent work of [2] shows that (in some sense) there can be a gap between these two quantities; there are cases where a learner provably cannot (i.e., in no Nash equilibrium of the game) achieve their expected Stackelberg value against a distribution of opponents.
@@
We crucially use the existence of efficient no-swap regret algorithms in constructing our optimal no-regret algorithm in Section 3. The first no-swap-regret algorithms were introduced by [16], who also demonstrated that their dynamics converge to correlated equilibria. Since then, numerous researchers have developed learning algorithms aimed at minimizing swap regret in games [20, 5, 29, 13].
@@ 3 Optimal No-Regret Commitment
Recall that we assume that the optimizer is one of $k$ different types, drawn from a known distribution $\mathcal{D}$ where type $i$ has probability $\alpha_{i}$ of occurring and corresponds to the payoff function $u_{O,i}$. Under the optimal menu $\mathcal{M}$, each of these different optimizers will pick some CSP in $\mathcal{M}$. We call the $k$-tuple of CSPs $\Phi=(\varphi_{1},\varphi_{2},\dots,\varphi_{k})\in\Delta_{mn}^{k}$ a valid CSP assignment of this menu $\mathcal{M}$ if $\varphi_{i}\in\arg max_{\varphi}u_{O,i}(\varphi)$ for all $i$. A given menu may have multiple valid CSP assignments due to ties. To resolve this, we assume that the optimizer breaks ties in favor of the learner ( any further tie-breaking rules does not affect the learner’s utility). However, we do not explicitly enforce this constraint; instead, by searching over all valid assignments, we inherently identify the optimal one from the perspective of the learner’s utility.
@@ 4.2 Hardness of testing the validity of menus
To prove Theorem 4.15, we will demonstrate how to reduce the approachability problem in Theorem 4.16 to an instance of our problem of determining whether a menu is valid.

### v2

@@ Abstract
We consider the problem of a learning agent who has to repeatedly play a general sum game against a strategic opponent who acts to maximize their own payoff by optimally responding against the learner’s algorithm. The learning agent knows their own payoff function, but is uncertain about the payoff of their opponent (knowing only that it is drawn from some distribution $\mathcal{D}$). What learning algorithm should the agent run in order to maximize their own total utility, either in expectation or in the worst-case over $\mathcal{D}$?
When the learning algorithm is constrained to be a no-regret algorithm, we demonstrate how to efficiently construct an optimal learning algorithm (asymptotically achieving the optimal utility) in polynomial time for both the in-expectation and worst-case problems, independent of any other assumptions. When the learning algorithm is not constrained to no-regret, we show how to construct an $\varepsilon$-optimal learning algorithm (obtaining average utility within $\varepsilon$ of the optimal utility) for both the in-expectation and worst-case problems in time polynomial in the size of the input and $1/\varepsilon$, when either the size of the game or the support of $\mathcal{D}$ is constant. Finally, for the special case of the maximin objective, where the learner wishes to maximize their minimum payoff over all possible optimizer types, we construct a learner algorithm that runs in polynomial time in each step and guarantees convergence to the optimal learner payoff. All of these results make use of recently developed machinery that converts the analysis of learning algorithms to the study of the class of corresponding geometric objects known as menus.
@@ 1.1 Our Results
    We then provide some evidence to show that the unconstrained problem above is hard when both $mn$ and $k$ are large. In particular, we prove that it is NP-hard to take as input a $k$-tuple $(u_{1},u_{2},\dots,u_{k})$ of real numbers and decide whether there exists a learning algorithm $\mathcal{A}$ where the optimizer of type $i$ receives utility at most $u_{i}$ when best responding to $\mathcal{A}$ (Theorem 4.15). This operation is a key subproblem in both of our approaches above, and implies that any efficient construction of an optimal learning algorithm must take a substantially different approach.
  4. 4.
    Finally, complementing this hardness result, we demonstrate an explicit efficient optimal learning algorithm under no additional constraints for the maximin objective, where the learner wants to maximize the utility they receive from the worst type of optimizer. Notably, this construction guarantees that the learner receives (nearly) their maximin value $V$ while sidestepping the actual problem of computing $V$ (a problem we conjecture is NP-hard).
@@
Finally, one may ask if it is truly necessary to pay this exponential cost in the dimension $\min(mn,k)$ when designing unconstrained learning algorithms. On one hand, this is a central roadblock to the Blackwell approachability based algorithms above – [29] show that the problem of deciding whether a specific set is NP-hard. Nonetheless, we show how to sidestep this obstacle for the specific objective of the maximin utility [^2] We briefly emphasize here that although we have stated the above results in the context of maximizing the expected utility of the learner when the optimizer is drawn from a known prior distribution, this optimization framework is fairly generic and allows for optimizing many other objectives over the space of learning algorithms (essentially, any objective which can be written as a concave function of the CSP assignment). of the learner (the worst-case utility against any optimizer in the support of $\mathcal{D}$). The key idea is to essentially run the same algorithm as before (computing a candidate optimal menu), but to initially be very optimistic about which menus are approachable. As we actually run our learning algorithm, either we will approach this menu without issue, or at some point our approachability algorithm will fail to play a valid action – at this point, we can restart the algorithm, being slightly more cautious about which menus are approachable. Interestingly, unlike all previous algorithms we have discussed, this algorithm does not allow us to directly compute the true optimal menu or even the learner’s optimal maximin value (in some sense, it relies on the optimizer to perform part of the computation). It is also essential to our approach that the maximin objective is in some sense “one-dimensional” – we leave it as an interesting open question whether this technique can be generalized to the original expected utility objective.
@@ 1.2 Related Work
There is also a rich line of work focused on the problem of learning the (single-shot) Stackelberg equilibrium of an unknown game via repeated interaction with an agent of some specified behavior. For example, one can try to estimate this Stackelberg equilibrium by repeatedly interacting with an agent that myopically best-responds every round [26, 3, 34], a no-regret agent [7, 20], a no-swap-regret agent / an agent that best-responds to calibrated action forecasts [7, 23], or an agent belonging to a parametrized class of behaviors [32]. It is worth emphasizing that this is a fairly different computational problem than the problem we study in this work – although the algorithms we commit to may “try” to learn the payoffs of their opponent in a similar way to these algorithms, we 1. do not make any behavioral assumptions about the opponents (allowing them to react fully rationally and non-myopically), and 2. attempt to find the Bayesian Stackelberg equilibrium of the full repeated game, not the single-shot Stackelberg equilibrium. Indeed, the recent work of [2] shows that (in some sense) there can be a gap between these two quantities; there are cases where a learner provably cannot (i.e., in no Nash equilibrium of the game) achieve their expected Stackelberg value against a distribution of opponents.
@@
We crucially use the existence of efficient no-swap regret algorithms in constructing our optimal no-regret algorithm in Section 3. The first no-swap-regret algorithms were introduced by [18], who also demonstrated that their dynamics converge to correlated equilibria. Since then, numerous researchers have developed learning algorithms aimed at minimizing swap regret in games [22, 5, 31, 13]. Our polynomial-time algorithm for designing general menus for the maximin objective is similar to (and partially motivated by) the technique of semi-separation recently introduced by [14], who use this technique to design learning algorithms that minimize a variant of swap regret (linear swap regret) despite this variant being computationally hard to compute.
@@ 3 Optimal No-Regret Commitment
Recall that we assume that the optimizer is one of $k$ different types, drawn from a known distribution $\mathcal{D}$ where type $i$ has probability $\alpha_{i}$ of occurring and corresponds to the payoff function $u_{O,i}$. Under the optimal menu $\mathcal{M}$, each of these different optimizers will pick some CSP in $\mathcal{M}$. We call the $k$-tuple of CSPs $\Phi=(\varphi_{1},\varphi_{2},\dots,\varphi_{k})\in\Delta_{mn}^{k}$ a valid CSP assignment of this menu $\mathcal{M}$ if $\varphi_{i}\in\arg\max_{\varphi}u_{O,i}(\varphi)$ for all $i$. A given menu may have multiple valid CSP assignments due to ties. To resolve this, we assume that the optimizer breaks ties in favor of the learner ( any further tie-breaking rules does not affect the learner’s utility). However, we do not explicitly enforce this constraint; instead, by searching over all valid assignments, we inherently identify the optimal one from the perspective of the learner’s utility.
@@ 4.2 Hardness of testing the validity of menus
To prove Theorem 4.15, we will demonstrate how to reduce the approachability problem in Theorem 4.16 to an instance of our problem of determining whether a menu is valid. The proof is deferred to Appendix A.
4.3 A Polynomial Time Algorithm for the Maximin Objective
In the previous section, we showed that it is computationally hard to determine whether a given menu is valid. If even checking whether a single menu is hard, this may suggest that it is hard to design a learning algorithm that implements the optimal menu.
Somewhat counterintuitively, in this section we will show that this is possible, at least in certain settings. In particular, we will design learning algorithms (without any additional constraints) that are nearly optimal for maximizing the learner’s maximin utility; the worst-case utility they might receive against any type of opponent.
For any $u_{L}$ and set of $k$ optimizers $\{u_{O,i}\}_{i=1}^{k}$, let $\textsf{OPT}=\sup_{\mathcal{M}}\min_{i\in[k]}V_{L}(\mathcal{M},u_{O,i})$ denote the best asymptotic per-round maximin utility a learner can achieve against this set of optimizers (where the supremum is over all valid menus $\mathcal{M}$). We have the following theorem.
Theorem 4.17.
For any fixed ${\varepsilon}>0$, there exists a learning algorithm $\mathcal{A}_{{\varepsilon}}$ (Algorithm 2) that guarantees a per-round maximin utility of
$\min_{i\in[k]}V_{L}(\mathcal{M}(\mathcal{A}_{{\varepsilon}}),u_{O,i})\geq\textsf{OPT}-{\varepsilon}$
and runs in time $\operatorname{poly}(n,m,k)$ per iteration.
The primary insight behind the design of the algorithm in Theorem 4.17 is that even if we do not know whether a specific menu is approachable (in the sense of Blackwell approachability, 4.7), we can still attempt to run an algorithmic instantiation of the Blackwell approachability theorem to attempt to approach this set.
What happens when we do try to approach an unapproachable set $\mathcal{M}$? Most algorithms for Blackwell approachability do not immediately fail when run on such sets (indeed, if they did, that would contradict the hardness result of [29]). Instead, they will approach $\mathcal{M}$ for some time until at some round (depending on the behavior of the adversary) they eventually fail to play an action, certifying that the set $\mathcal{M}$ is unapproachable.
We demonstrate this for the specific application of Blackwell Approachability we need in Theorem 4.17. Consider the approachability game defined earlier where $u:\Delta_{m}\times\Delta_{n}\rightarrow\mathbb{R}^{mn}$ is defined via $u(x,y)=x\otimes y$, and let $\mathcal{C}(\Phi)$ be the candidate menu for an arbitrary incentive-compatible CSP assignment $\Phi=(\varphi_{1},\varphi_{2},\dots,\varphi_{k})\in\mathcal{R}$. We will try to force $u$ to approach $\mathcal{C}(\Phi)$. Note that while we assume $\Phi$ is incentive-compatible, it is not necessarily the case that $\Phi\in\mathcal{S}$, and so $\mathcal{C}(\Phi)$ is not necessarily response satisfiable with respect to $u$.
We say that a learning algorithm is abortable if the learning algorithm has the option to terminate its own execution after a finite number of rounds (i.e., outputting an Abort signal instead of playing an action) [^3] We include a formal definition of what this means (and how this affects the definition of menus) in Appendix B.4.. The following theorem provides guarantees on an abortable learning algorithm for approaching the set $\mathcal{C}(\Phi)$ for any incentive-compatible $\Phi$ (in particular, note that the quantity on the LHS of (8) can be thought of as the maximum extent to which the CSP $\varphi=(1/T)\cdot(\sum_{t}x_{t}\otimes y_{t})$ violates the constraints of $\mathcal{C}(\Phi)$).
Theorem 4.18.
For any incentive-compatible CSP assignment $\Phi\in\mathcal{R}$, there is an efficient abortable learning algorithm $\textsc{BlackwellAbort}(\Phi)$ with the following properties.
  • If at round $\tau$, $\textsc{BlackwellAbort}(\Phi)$ has not stopped, then its transcript of play up to this round satisfies:
    $\max_{i\in[k]}\left(\sum_{t=1}^{\tau}(u_{O,i}(x_{t},y_{t})-u_{O,i}(\varphi_{i}))\right)=O(\sqrt{\tau\log k}).$ (8)
    In particular, the menu $\mathcal{M}(\textsc{BlackwellAbort}(\Phi))$ is contained within the set $\mathcal{C}(\Phi)$.
  • If the learning algorithm $\textsc{BlackwellAbort}(\Phi)$ ever aborts, then the set $\mathcal{C}(\Phi)$ is not response-satisfiable with respect to $u$.
  • Each $\varphi_{i}\in\Phi$ is contained in the menu $\mathcal{M}(\textsc{BlackwellAbort}(\Phi))$; i.e., there is a sequence of opponent actions against which this algorithm will never abort and where the CSP will converge to $\varphi_{i}$.
Proof.
Consider the following learning algorithm:
  • Initialize an instance of Anytime Hedge [17] with $k$ actions (this algorithm has the guarantee that the external regret at round $t$ is at most $O(\sqrt{t\log k})$ for all $t\geq 1$).
  • For each round $t$ (for $t\geq 1$), do the following:
    • Receive the action $p^{t}\in\Delta_{k}$ from the instance of Anytime Hedge.
    • Choose an $x_{t}\in\Delta_{m}$ s.t.
      $\sum_{i=1}^{k}p^{t}_{i}(u_{O,i}(x_{t},y)-u_{O,i}(\varphi_{i}))\leq 0$
      for all $y\in\Delta_{n}$. We can find such an $x_{t}$ by solving a linear program; if no such $x_{t}$ exists, abort.
    • Play this $x_{t}$ and observe the opponents action $y_{t}\in\Delta_{n}$
    • Feed Hedge the reward vector $r^{t}\in[-1,1]^{k}$ defined via $r^{t}_{i}=u_{O,i}(x_{t},y_{t})-u_{O,i}(\varphi_{i})$.
We claim that this learning algorithm has the desired guarantees. To see this, first note that if the above algorithm ever aborts, by von Neumann’s minimax theorem, it must be the case that there exists a $y^{*}\in\Delta_{n}$ such that
$\sum_{i=1}^{k}p^{t}_{i}(u_{O,i}(x,y^{*})-u_{O,i}(\varphi_{i}))>0$
for all $x\in\Delta_{m}$. But now, $y^{*}$ is a certificate that $\mathcal{C}(\Phi)$ is not response-satisfiable with respect to $u$ (i.e., there is no $x$ such that $u(x,y^{*})=x\otimes y^{*}$ satisfies all the constraints of $\mathcal{C}(\Phi)$).
On the other hand, note that if the algorithm has successfully run for $\tau$ rounds, then
$\max_{i\in[k]}\left(\sum_{t=1}^{\tau}(u_{O,i}(x_{t},y_{t})-u_{O,i}(\varphi_{i}))\right)$
$\leq$ $\max_{i\in[k]}\left(\sum_{t=1}^{\tau}(u_{O,i}(x_{t},y_{t})-u_{O,i}(\varphi_{i}))\right)-\sum_{t=1}^{\tau}\sum_{i=1}^{k}p^{t}_{i}(u_{O,i}(x,y^{*})-u_{O,i}(\varphi_{i}))$
$=$ $\max_{i\in[k]}\sum_{t=1}^{\tau}r_{t,i}-\sum_{t=1}^{\tau}\langle p^{t},r_{t}\rangle$
$=$ $\operatorname{\mathsf{Reg}}(\mathrm{AnytimeHedge})=O(\sqrt{\tau\log k}).$
Finally, this algorithm, as written, does not guarantee that $\varphi_{i}$ belongs to its menu. However, we can easily modify this algorithm in the same way as Lemma 3.5 of [1] by adding a cheap talk phase (or implementing one through the details of the protocol) where we cooperate with the opponent to play a sequence of actions converging to the $\varphi_{i}$ of their choice. See also the discussion in Section B.3 for more details. ∎
Algorithm 2 Polynomial Time Learning Algorithm for Maximin Value
1:  Input: parameter ${\varepsilon}>0$
2:  Initialize $V=1$An upper bound on OPT}
3:  while true do
4:     Set $\mathcal{F}(V)\leftarrow\{\phi\mid u_{L}(\phi)\geq V\}$
5:     Set CSP assignment $\Phi(V)\leftarrow\{\phi_{i}\}_{i\in[k]}$ where:
6:         $\phi_{i}\in\operatorname*{arg\,max}_{\phi\in\mathcal{F}(V)}u_{O,i}(\phi)$
7:     {Note that $\Phi(V)$ is incentive-compatible ($\Phi(V)\in\mathcal{R}$) since each $\varphi_{i}$ maximizes $u_{O,i}$ over all CSPs in $\mathcal{F}(V)$.}
8:     Run the learning algorithm $\textsc{BlackwellAbort}(\Phi(V))$ (see Theorem 4.18) until it aborts.
9:     Set $V\leftarrow V-{\varepsilon}$note that this is only reached if the previous step aborts}
10:  end while
Based on this idea, we show how to construct an efficient algorithm that maximizes the maximin objective in the following way (described in Algorithm 2). The core idea is to maintain an upper bound $V$ on OPT. The algorithm runs in epochs, where in each epoch we find the opponent-optimal CSP assignment $\Phi$ supported on CSPs that guarantee the learner at least utility $V$. We then try to approach $\mathcal{C}(\Phi)$ by running the abortable learning algorithm in Theorem 4.18. If we abort, we decrease $V$ a little bit and start a new epoch.
We can now prove Thorem 4.17.
@@ 4.3 A Polynomial Time Algorithm for the Maximin Objective
We first make the following observation: at every step of the game, the variable $V$ in Algorithm 2 never drops below $\textsf{OPT}-{\varepsilon}$. To see this, assume $V\leq\textsf{OPT}$, and consider the CSP assignment $\Phi$ constructed by picking $\varphi_{i}\in\operatorname*{arg\,max}_{\varphi\in\mathcal{F}(V)}u_{O,i}(\varphi)$. We will show that $\mathcal{C}(\Phi)$ is approachable, which implies that the learning algorithm $\mathcal{A}$ in Theorem 4.18 will never abort when run on $\Phi$ (and therefore, $V$ never further decreases after the first time it drops below OPT).
Consider the CSP assignment $\Phi^{\prime}$ constructed from $\mathcal{F}(\textsf{OPT})$. Note that $\mathcal{C}(\Phi^{\prime})\subseteq\mathcal{C}(\Phi)$; this follows almost directly from the fact that $\mathcal{F}(\textsf{OPT})\subseteq\mathcal{F}(V)$, implying $u_{O,i}(\varphi)\geq u_{O,i}(\varphi^{\prime})$, which in turn guarantees that each constraint hyperplane of $\mathcal{C}(\varphi)$ is weaker than the corresponding constraint in $\mathcal{C}(\varphi^{\prime})$. Now, we also know that an optimal menu $\mathcal{M}$ exists and attains OPT – let $\Phi^{*}$ be the CSP assignment (i.e. $\varphi^{*}_{i}$ is the profile in $\mathcal{M}$ maximizing the optimizer type’s utility). Clearly, $u_{O,i}(\varphi^{*}_{i})\leq u_{O,i}(\varphi^{\prime}_{i})$, which directly implies that $\mathcal{C}(\varphi^{*})\subseteq\mathcal{C}(\varphi^{\prime})$. But, tautologically, $\mathcal{M}\subseteq\mathcal{C}(\Phi^{*})$ and therefore, since approachability is upwards closed, $\mathcal{C}(\Phi)$ is approachable.
Armed with this observation, we now examine the menu $\mathcal{M}(\mathcal{A}_{{\varepsilon}})$ of our algorithm. Let $\mathcal{V}=\{1,1-{\varepsilon},1-2{\varepsilon},\dots,V_{\min}\}$ be the set of possible values that any adversary can cause $V$ to take on; note that it contains every real number of the form $1-n{\varepsilon}$ in the range $(\textsf{OPT}-{\varepsilon},1]$ (in particular, conversely to the above observation, an adversary can always cause $V$ to decrease if $V>\textsf{OPT}$ by repeatedly playing a $y\in\Delta_{n}$ that demonstrates that $\mathcal{C}(V)$ is not approachable). We claim that $\mathcal{M}(\mathcal{A}_{{\varepsilon}})$ has the following properties:
  1. 1.
    For each $V\in\mathcal{V}$, every CSP $\varphi$ in the CSP assignment $\Phi(V)$ belongs to $\mathcal{M}(\mathcal{A}_{{\varepsilon}})$.
  2. 2.
    The menu $\mathcal{M}(\mathcal{A}_{{\varepsilon}})$ is a subset of $\mathcal{C}(\Phi(V_{\min}))$.
The first property is true since each such CSP is explicitly added to one of the menu of one of the potential instances of BlackwellAbort. The second property is true [^4] Here we are implicitly using the fact that the menu of an algorithm constructed in this way is a subset of the convex hull of all of the abortable sub-algorithms. See Lemma B.2 in Appendix B.4 for a proof of this. because each of the individual menus $\mathcal{M}(\textsc{BlackwellAbort}(\Phi(V))$ is contained within $\mathcal{C}(\Phi(V))$ (by the regret guarantee in Theorem 4.18), and these menus are nested ($\mathcal{C}(\Phi(1))\subseteq\mathcal{C}(\Phi(1-{\varepsilon}))\subseteq\dots\subseteq\mathcal{C}(\Phi(V_{\min}))$). Note that every $\varphi_{i}\in\Phi(V)$ is in fact contained in the set $\mathcal{C}(\Phi(V))$, i.e., $\Phi(V)$ satisfies the incentive compatibility conditions ($\Phi(V)\in\mathcal{R}$) – this is because each $\varphi_{i}$ is the maximum value point in $F(V)$ for optimizer payoff $u_{O,i}$.
We can now evaluate the values $V_{L}(\mathcal{M}(\mathcal{A}_{{\varepsilon}}),u_{O,i})$. Consider the CSP $\varphi_{i}$ that the optimizer $u_{O,i}$ picks from $\mathcal{M}(\mathcal{A}_{{\varepsilon}})$. Let $\varphi^{\prime}_{i}$ be this optimizer’s CSP assignment in $\Phi(V_{\min})$. We claim that $u_{O,i}(\varphi^{\prime}_{i})=u_{O,i}(\varphi_{i})$. Indeed, this follows from the fact that $\mathcal{M}(\mathcal{A}_{{\varepsilon}})\subseteq\mathcal{C}(\Phi(V_{\min}))$, and by definition no CSP in $\mathcal{C}(\Phi(V_{\min}))$ has better utility for optimizer $i$ than their assigned one. On the other hand, we also know that $\varphi^{\prime}_{i}\in\mathcal{M}(\mathcal{A}_{{\varepsilon}})$. So, since the optimizer breaks ties in favor of the learner, we must have $V_{L}(\mathcal{M}(\mathcal{A}_{{\varepsilon}}),u_{O,i})\geq u_{L}(\varphi^{\prime}_{i})$. But finally, since $\varphi^{\prime}_{i}$ was chosen from $\mathcal{F}(V_{\min})$, it must be the case that $u_{L}(\varphi^{\prime}_{i})\geq V_{\min}\geq\textsf{OPT}-{\varepsilon}$, as desired. ∎
It is interesting to consider the strategic situation faced by one of the opponents facing Algorithm 2. We can think of each stage of Algorithm 2 as a challenge from the learner to the opponent of the form: “I think I can keep the average CSP in the set $\mathcal{C}(\Phi(V))$. If I can, it is best for you to play your assigned CSP $\varphi_{i}\in\Phi$ (and I will cooperate with you to reach that CSP). Otherwise, you must demonstrate to me that the set $\mathcal{C}(\Phi(V))$ is in fact not approachable (at which point I will revise my play).” In this way, the learner delegates the computational complexity of computing the optimal menu $\mathcal{M}$ to their opponent, who either must help the learner compute this menu or settle for a potentially worse outcome. As a result, although the guarantee of Theorem 4.17 is stated only in terms of the guarantee achievable against a perfectly rational (and computationally unconstrained) opponent, it also makes sense (and in fact, is even better) to play Algorithm 2 against computationally-bounded rational opponents.
Finally, in this section we have restricted ourself to the maximin objective. It is natural to wonder whether we can achieve a similar guarantee for the expected utility objective studied throughout the rest of the paper – we leave this as a major open question. We note that the current approach takes advantage in several places of the “one-dimensional” nature of the maximin objective (for example, it is important that the optimal CSP assignment is simply the one that assigns each opponent their favorite CSP with value at least $V$ for the optimizer), and it is not clear how to adapt it to these more general objectives.